40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jul 31 21:12:05 2013, manasa, Update, General, Pump down called off 
    Reply  Thu Aug 1 16:03:52 2013, Steve, Update, General, reopen BS & IOO IOOnorth.jpg
    Reply  Thu Aug 1 22:49:31 2013, manasa, Update, General, Green status after PR3 flipping investigated 
       Reply  Fri Aug 2 17:50:10 2013, Jenne, Update, General, PR3 wedge angle adjusted 
          Reply  Fri Aug 2 22:51:10 2013, Jenne, Update, General, All vent tasks complete, just need oplev check 
             Reply  Mon Aug 5 18:02:34 2013, manasa, Update, General, Close up list checked 
             Reply  Tue Aug 6 07:52:58 2013, manasa, Update, General, Ready to pump down MCspot_0806.pngOSEM_0806.pngOL_0806.pngQPD_0806.png
Message ID: 8962     Entry time: Fri Aug 2 22:51:10 2013     In reply to: 8960     Reply to this: 8965   8969
Author: Jenne 
Type: Update 
Category: General 
Subject: All vent tasks complete, just need oplev check 

[Manasa, Koji, Jenne]

We went into the BS and IOO chambers, and aligned the green beams such that they came out of the vacuum chamber.  The idea here was to get the beams at the same height, but slightly offset in yaw.  This required moving the Periscope on BS table, PBS in front of that periscope, the Periscope on the IOO table, and 2 steering mirrors on the IOO table after the 2nd periscope.  The tables were not releveled, although we have aligned the full interferometer to this situation, so we do not want to touch the tables.  The MC spot positions are still consistent with those measured earlier this afternoon, before this work, so I'm not concerned.

We confirmed that both green beams are hitting a good place (centered in pitch, and just left and right of center in yaw) on the mirror in the OMC chamber, and are getting to the center of the first mirror on the PSL table.  We then coarsely aligned the beams on the PSL table.

We then relocked and aligned the arms for IR, and checked that the AS beam is centered on the mirrors in the BS chamber, and that the beam is coming out, and to the AS table.  I touched the last mirror before PZT3 a small amount in yaw, and then PZT3 in pitch and yaw, until we saw the beam recentered on the first mirror on the AS table.  At that point, we were also back to the center of the AS camera (which is good, since Koji had aligned all of that the other day).  So, the AS beam is good.

We checked IPPOS, and have centered the beam on all the mirrors, and aligned the beam onto the QPD. 

We checked IPANG, by looking through the viewports at the mirrors in the ETMY chamber.  We are now centered in yaw, but clipping a bit low.  This is what we want, since we always end up drifting high during the pump-down. 

We see a nice, unclipped REFL beam on the camera.

We see a beam a little high on the POP camera, but Koji looked on the table with a card, and saw the beam....we just need to do minor alignment on the out of vac mirrors.

We checked again that the green TEM00 beams from both arms come to the PSL table. 

We are getting POX and POY out, since we are using them to lock and align the arms.

Manasa and Koji recovered one clean allen key from the bottom of the chambers, but one remains, as a sacrifice to the vacuum gods. 

I believe that, with the exception of checking the oplevs and taking photos of PR3, and the green steering optics, we have finished all of our vent tasks.  We should do a quickie alignment on Monday, check the oplevs, take some photos, and put on the heavy doors.  Pumping can start either Monday afternoon or Tuesday morning. 

 

ELOG V3.1.3-